Suppose that the utility function for two commodities is:

U(q1, q2) = q1α q2(1-α)

Let the prices of the two commodities be p1 and p2 and let the consumer’s income be M.

(1) Check the properties of marginal utilities. In particular, check whether it satisfies diminishing marginal utilities.

(2) Assuming all income is spent on these two commodities, derive the demand curves for the two commodities.

(3) What happens if U(q1, q2) = q1α q2β?


1
Expert's answer
2021-08-02T08:33:36-0400

marginal utility(MU) is the addition to the total utility when an additional unit of a good is consumed. The law of diminishing marginal utility states that when more and more units of one good are consumed, the marginal utility from each successive unit declines. 

The marginal utility of good q1 is the derivative of total utility with respect to q1. The marginal utility of good q2 is the derivative of total utility with respect to q2.


"1)\\
U(q_1,q_2) =q_1^αq_2^{(1-α)}\\MUq_1=\frac{∂U(q_1,q_2)}{∂q_1}\\=α\frac{q_2^{(1-α)}}{q_1^{(1-α)}}\\MUq_2=\frac{∂U(q_1,q_2)}{∂q_2}\\=(1-α)\frac{q_1^α}{q_2^α}"

The marginal utility is diminishing because the quantity is in the denominator. So, the marginal utility of q1 decreases with the increase in q1. The marginal utility of q2 decreases with the increase in q2


2)

A consumer consumes where its utility is maximized. The consumer maximizes its utility where the marginal rate of substitution(MRS) is equal to the price ratio. The marginal rate of substitution is the ratio of marginal utility of good 1 to the marginal utility of good 2. 


"MRS=\frac{\frac{αq_2^{(1−α)}}{q_1^{(1−α)}}}{\frac{(1−α)q_1^α}{q_2^α}}"

"=\frac{αq_2}{(1−α)q_1}"


"MRS=\frac{p_1}{P_2}\\\frac{αq_2}{(1−α)q_1}=\frac{p_1}{P_2}"

"q_2=\frac{(1−α)q_1p_1}{αp_2}"

Budget constraint

"p_1q_1 +p_2q_2=M"

"p_1q_1 +p_2(\frac{(1−α)q_1p_1}{αp_2})=M"

"p_1q_1 +\frac{(1−α)q_1p_1}{α}=M"

"p_1q_1 \frac{(1+1−α)}{α}=M"

"q_1=\frac{α×M}{p_1}" ...demand function for q1

"q_2=\frac{(1−α)p_1(\frac{α×M}{p_1})}{αp_2}"

"=\frac{(1−α)×M}{p_2}" ...demand function for q2


3)

As the power of q2 changes from "1-α\space to\space β" , the demand function of q2. The demand function of q1 remains the same.

"q_2 =\frac{β×I}{p_2}"